OpenStudy (anonymous):

physics question @ganeshie8 @nincompoop @wio

ganeshie8 (ganeshie8):

I think you just need to split the forces into components and solve : \[-T_1 \sin(65)+T_2\sin(80) -900\sin(\alpha-90)=0\tag{1}\] \[T_1 \cos(65)+T_2\cos(80) -900\cos(\alpha-90)=0\tag{2}\]